OpenStudy (anonymous):

express the product as a radical expression (x^3)^3/4

OpenStudy (swissgirl):

umm do u know the laws of exponents? \( (x^m)^n = x^{\large m*n} \)

OpenStudy (swissgirl):

ohhhh i misread it lol it says \( \huge (x^3)^{ 3 \over 4} \)

OpenStudy (swissgirl):

Ok so the lasw of exponents r as follows: \( \huge x^{ m\over n} = \sqrt[n]{x^m} \)

OpenStudy (anonymous):

okayyy(:

OpenStudy (swissgirl):

ok so here it wld be: \( \large \sqrt[4]{(x^3)^3} =\sqrt[4]{(x^9)}\)

OpenStudy (swissgirl):

which can actually be reduced to: \( \large x^2 \sqrt[4]{x} \)

OpenStudy (anonymous):

thank you sooo much you are great!:D

OpenStudy (swissgirl):

yup sorry abt the mistake b4 this is the correct answer

OpenStudy (anonymous):

thnx so much and it was no biggie:)